OpenStudy (anonymous):

Simplify. square root of 9 over sixth root of 9

OpenStudy (marvin456):

21? @Vniverse

OpenStudy (anonymous):

uhhh 21

OpenStudy (anonymous):

9 to the power of negative 2 over 3 9 to the power of negative 1 over 3 8 to the power of 3 over 10 9 to the power of 1 over 3

OpenStudy (anonymous):

These are the options. I don't understand radicals at all though

OpenStudy (anonymous):

So, D?

OpenStudy (anonymous):

9^(1/2)/9^(1/6) = 9^(1/2-1/6) = 9^(2/3)

OpenStudy (anonymous):

So, A? Right?

OpenStudy (anonymous):

They're all wrong

OpenStudy (anonymous):

How?